# Point Linguist Model: Segment Any Object via Bridged Large 3D-Language Model

### v1

@@ IV-C Per-class results
We present per-class results on the OVIS task in Table VII and Table VIII. The results are reported in terms of AP50, with performance on novel classes highlighted in teal. On the ScanNetV2 dataset, PLM consistently outperforms OpenIns3D across most novel classes, demonstrating its superior generalization ability in the open-vocabulary setting. Compared to PLA [52], PLM achieves more stable and higher recognition rates for underrepresented categories, especially in challenging categories like picture (41.2%), counter (20.4%), and desk (16.9%). PLM outperforms OpenIns3D in key furniture categories, such as cabinet (30.2% vs. 24.3%) and bathtub (64.3% vs. 7.6%), indicating its stronger ability to segment complex object classes. On the S3DIS dataset, PLM achieves competitive performance in architectural components and furniture classes, such as window (40.7%), door (52.8%), table (25.6%), and board (57.0%). Compared to OpenIns3D, PLM improves segmentation performance on chairs (67.6% vs. 52.0%) and bookcases (25.9% vs. 0.0%), demonstrating strong semantic transferability in the open-vocabulary setting.
@@ IV-D Additional Visualization
V Conclusion

### v2

@@ IV-C Per-class results
We present per-class results on the OVIS task in Table XI and Table XII. The results are reported in terms of AP50, with performance on novel classes highlighted in teal. On the ScanNetV2 dataset, PLM consistently outperforms OpenIns3D across most novel classes, demonstrating its superior generalization ability in the open-vocabulary setting. Compared to PLA [10], PLM achieves more stable and higher recognition rates for underrepresented categories, especially in challenging categories like picture (41.2%), counter (20.4%), and desk (16.9%). PLM outperforms OpenIns3D in key furniture categories, such as cabinet (30.2% vs. 24.3%) and bathtub (64.3% vs. 7.6%), indicating its stronger ability to segment complex object classes. On the S3DIS dataset, PLM achieves competitive performance in architectural components and furniture classes, such as window (40.7%), door (52.8%), table (25.6%), and board (57.0%). Compared to OpenIns3D, PLM improves segmentation performance on chairs (67.6% vs. 52.0%) and bookcases (25.9% vs. 0.0%), demonstrating strong semantic transferability in the open-vocabulary setting.
IV-D Computational Performance of the PLM
For modern multimodal LLM-based methods, the LLM component typically dominates the overall computational cost. Since different approaches may adopt different LLM backbones, computational comparisons across methods are often less indicative of the efficiency of the proposed 3D reasoning modules themselves.
Under these constraints, we aim to provide meaningful efficiency analysis from multiple complementary perspectives. First, to offer a coarse latency reference, we compare our inference latency with representative non-LLM-based methods, including OneFormer3D [29] (221.00 ms), UniSeg3D [63] (230.03 ms), SSTNet [33] (400 ms), and PointGroup [26] (372 ms). While these methods are not directly comparable due to architectural differences, the comparison provides a useful sense of the practical runtime scale.
We then report a detailed computational analysis of our model in Table XIII. Specifically, we report forward, backward, and total computation (in TFLOPs), throughput in samples per second, per-sample latency, as well as GPU memory usage. All results are measured using a batch size of 12 under the same hardware configuration. Under the default setting with GRD Decoder depth nn=3, our model achieves a throughput of 3.2 samples/sec with a latency of 312.76 ms. To further examine the computational efficiency of our design, we analyze the impact of increasing the GRD depth. As shown in Table XIII, increasing the GRD depth from nn=3 to nn=5 results in only a marginal increase in total computation (141.56 → 146.36 TFLOPs), while the throughput (3.2 → 3.1 samples/sec) and latency (312.76 → 326.79 ms) remain largely unchanged.
V Conclusion
-A Additional Loss Details
Notation
  • $M$: Number of predicted masks.
  • NN: Number of points per mask.
  • $\mathbf{y}=\{\mathbf{y}_{j}\}_{j=1}^{M}$: Set of ground-truth masks (including both targets and distractors), where each $\mathbf{y}_{j}$ comprises a binary mask and a class label.
  • $\hat{\mathbf{y}}=\{\hat{\mathbf{y}}_{i}\}_{i=1}^{M}$: Set of predictions, where
    $\hat{\mathbf{y}}_{i}=\bigl\{\mathbf{m}_{i},\;\mathbf{c}_{i},\;\mathbf{b}_{i}\bigr\},$
    with
    • $\mathbf{m}_{i}=(p_{i}^{1},\dots,p_{i}^{N})\in[0,1]^{N}$: predicted mask probability vector,
    • $\mathbf{c}_{i}=[p_{i,0},\,p_{i,1}]^{\top}\in[0,1]^{2}$: predicted class probabilities (0 = target, $1$ = distractor),
    • $\mathbf{b}_{i}\in\mathbb{R}^{6}$: predicted bounding box, parameterized by center $(x,y,z)\in\mathbb{R}^{3}$ and size $(w,h,d)\in\mathbb{R}^{3}$.
  • $\mathbf{y}_{\text{mask}}=(g^{1},\dots,g^{N})\in\{0,1\}^{N}$: ground-truth mask vector.
  • $\mathbf{y}_{\text{cls}}=[y_{0},y_{1}]^{\top}\in\{0,1\}^{2}$: one-hot class label.
  • $B_{i}$ and $B_{\text{gt}}$: volumes of the predicted and ground-truth boxes, respectively; $C$ is the smallest enclosing box of $B_{i}\cup B_{\text{gt}}$, with volume $|C|$.
  • Hyperparameters $\lambda_{\text{cls}},\;\lambda_{\text{mask}},\;\lambda_{\text{box}}$: weights for the classification, mask, and box losses.
Hungarian Matching
We perform one-to-one matching between $\hat{\mathbf{y}}$ and $\mathbf{y}$ via the Hungarian algorithm:
$\begin{split}\hat{\mathbf{y}}_{\text{pos}}&=\arg\min\;\mathfrak{L}_{\text{match}}\bigl(\hat{\mathbf{y}}_{i},\mathbf{y}\bigr),\\ \mathfrak{L}_{\text{match}}&=\lambda_{\text{cls}}\,\mathfrak{L}_{\text{cls}}+\lambda_{\text{mask}}\,\mathfrak{L}_{\text{mask}}.\end{split}$
Classification Loss (Focal Loss)
For each matched prediction $\hat{\mathbf{y}}_{i}$, we compute
$\mathfrak{L}_{\text{cls}}(\hat{\mathbf{y}}_{i},\mathbf{y}_{\text{cls}})=-\sum_{c\in\{0,1\}}\alpha_{c}\,(1-p_{i,c})^{\gamma}\,y_{c}\,\log(p_{i,c}),$
with $\alpha_{c}=1$ and $\gamma=1$ by default.
Mask Loss
The mask loss combines Dice loss and binary mask focal loss:
$\mathfrak{L}_{\text{mask}}=\mathfrak{L}_{\text{dice}}+\mathfrak{L}_{\text{bfocal}}.$
Dice Loss
$\mathfrak{L}_{\text{dice}}(\mathbf{m}_{i},\mathbf{y}_{\text{mask}})=1-\frac{2\sum_{n=1}^{N}p_{i}^{n}\,g^{n}}{\sum_{n=1}^{N}p_{i}^{n}+\sum_{n=1}^{N}g^{n}}\,.$
Binary Mask Focal Loss
$\begin{split}\mathfrak{L}_{\text{bfocal}}(\mathbf{m}_{i},\mathbf{y}_{\text{mask}})&=-\frac{1}{N}\sum_{n=1}^{N}\Bigl[(1-p_{i}^{n})^{\gamma}g^{n}\log(p_{i}^{n})\\ &\quad+(p_{i}^{n})^{\gamma}(1-g^{n})\log(1-p_{i}^{n})\Bigr],\end{split}$
with $\gamma=1$.
Box Loss
The box loss is the sum of an $L1$ loss and a Generalized IoU loss:
$\mathfrak{L}_{\text{box}}=\mathfrak{L}_{L1}+\mathfrak{L}_{\text{giou}}.$
$L1$ Loss
$\mathfrak{L}_{L1}(\mathbf{b}_{i},\mathbf{b}_{\text{gt}})=\bigl\lVert\mathbf{b}_{i}-\mathbf{b}_{\text{gt}}\bigr\rVert_{1}.$
Generalized IoU Loss
First define
$\begin{split}\mathrm{IoU}(B_{i},B_{\text{gt}})&=\frac{|B_{i}\cap B_{\text{gt}}|}{|B_{i}\cup B_{\text{gt}}|},\\ \mathrm{GIoU}(B_{i},B_{\text{gt}})&=\mathrm{IoU}(B_{i},B_{\text{gt}})-\frac{|C\setminus(B_{i}\cup B_{\text{gt}})|}{|C|},\end{split}$
then
$\mathfrak{L}_{\text{giou}}=1-\mathrm{GIoU}(B_{i},B_{\text{gt}}).$
Overall Loss
Only the matched prediction $\hat{\mathbf{y}}_{\text{pos}}$ contributes:
$\mathfrak{L}=\underbrace{\lambda_{\text{cls}}\,\mathfrak{L}_{\text{cls}}+\lambda_{\text{mask}}\,\mathfrak{L}_{\text{mask}}}_{\mathfrak{L}_{\text{match}}}+\lambda_{\text{box}}\,\mathfrak{L}_{\text{box}}.$
@@ V Conclusion
-C Limitation Discussion
The proposed OcDR and GRD are not limited to indoor scenes. In principle, our PLM can be applied to outdoor LiDAR data, CAD scenes, and settings with larger point counts. However, transfer to such domains is mainly constrained by the encoder rather than the PLM itself. Due to the substantial differences in scene scale, point density, and task formulation, effective transfer typically requires a well-aligned point–image–text encoder, which in turn depends on large-scale, high-quality multimodal paired data. Compared to indoor scenarios, such data remain relatively limited for outdoor LiDAR and CAD settings.